# Expanding the Generative AI Design Space through Structured Prompting and Multimodal Interfaces

### v1

@@ Abstract.
Text-based prompting remains the dominant interaction paradigm in generative AI, yet it often results in a high-friction experience for novice users, such as small business owners (SBOs), attempting to articulate creative or domain-specific goals for advertising. To investigate this challenge, we conducted a study with six SBOs in the United Kingdom, focusing on their advertising practices and perceptions and usage of AI tools in this context. Our findings surfaced two persistent breakdowns in current generative AI systems: first, the cognitive burden of prompt engineering, as users struggled to translate abstract creative goals into effective textual inputs; and second, the frequent generation of generic outputs that failed to align with users’ articulated brand vision. To address these issues, we developed ACAI (AI Co-Creation for Advertising and Inspiration), a multimodal, GenAI-powered advertisement creation tool designed to support novice designers by reimagining the prompt interface. ACAI features a structured, panel-based interface composed of three modules: the Branding Panel, the Audience & Goals Panel, and the Inspiration Board Panel to provide SBOs with outputs that align with their creative vision by reducing prompt ambiguity. This work contributes to HCI research on generative systems by showing how structured interfaces can foreground user-defined context to improve both alignment and promptability in novice workflows.
@@ 1. Introduction
To investigate how current generative systems support SBOs’ creative workflows, we conducted a formative study with six SBOs in Manchester, United Kingdom. Using semi-structured interviews, we examined participants’ advertising practices, experiences with generative AI, and expectations for AI-assisted design. To investigate how current generative systems support SBOs’ creative workflows, we conducted a formative study with six small business owners (SBOs) in Manchester, United Kingdom. Using semi-structured interviews, we examined participants’ advertising practices, experiences with generative AI, and expectations for AI-assisted design. Our findings surfaced three recurring challenges: (1) the cognitive effort required to translate abstract, domain-specific goals into usable prompts; (2) dissatisfaction with generic outputs that failed to reflect brand tone, values, or audience needs; and (3) concerns about losing creative control due to over-automation, signaling the need for systems that enable co-creation and preserve user agency. These themes align with ongoing HCI work on promptability (Morris, 2024) and the “gulf of envisioning” (Subramonyam et al., 2024), which highlight how non-expert users struggle to effectively guide generative systems.
@@
This work makes three primary contributions to HCI research on GenAI:
@@ 2.1. Contextualization in Generative and Interactive Systems
However, current text-to-image generation models primarily rely on text-based prompts and offer limited affordances for encoding structured user intent or domain-specific constraints (Lee et al., 2023). While these systems produce compelling outputs, they often fall short in supporting intention-aligned co-creation, particularly for novice users in task-specific domains. Motivated by these limitations, we position ACAI to bridge the gap between generic text-to-image systems and context-sensitive creative workflows by foregrounding explicit, user-defined contextualization.
@@ 3.1. Key Findings from Formative Research
We analyzed the transcripts using reflexive thematic analysis. The following three findings emerged, informing the design of ACAI’s interface and interaction model:
@@
    Challenges in Articulating Brand Intuition through Prompts: Participants struggled to articulate their brand vision using traditional text-based prompts. While they possessed a strong intuitive understanding of their brand identity, they found it difficult to externalize this understanding. One participant shared, “I instructed it to be funny, but that’s quite subjective, so I had to edit it quite heavily.” -P2. This finding underscores the tacit and iterative nature of branding, highlighting how limitations in promptability introduce friction into the creative workflows of novice users.
@@
    Tensions Between Generativity and Control: While participants valued the efficiency offered by generative AI, they expressed concerns about losing creative control. Many voiced discomfort with over-automation and emphasized a preference for systems that facilitate co-creation, offering opportunities to adjust or override AI-generated outputs. One participant explained, “You couldn’t really change it, when AI produced that for you, there weren’t tweaks you could make to it… so that’s the limitation of it.” (P3). These tensions underscore the importance of interaction models that preserve user agency within AI-augmented workflows.
@@
    Generic and Misaligned AI Outputs: Participants frequently described AI-generated advertisements as generic, repetitive, or misaligned with their brand’s tone and aesthetic. They noted that outputs often lacked the nuance necessary to reflect distinctive brand identities. This disconnect reinforces the need for generative systems that incorporate user-defined brand elements and contextual cues to produce brand-consistent outputs.
@@ 4. ACAI: AI Co-Creation for Advertising and Inspiration
Drawing on our formative study with small business owners (SBOs), we developed ACAI (AI Co-Creation for Advertising and Inspiration), a generative AI-powered prototype to support the creation of brand-aligned advertisements. ACAI is designed for users with limited design expertise but strong knowledge of their brand.
@@ 4.1. System Overview
ACAI was developed as a standalone prototype to facilitate user testing sessions focused on co-creation practices among novice designers. The system leverages Gemini 1.5 Pro, a multimodal large language model (MLLM), to support its generative capabilities. ACAI’s architecture is organized into three core components: the User Input Layer, the Prompt Processing Layer, and the Output Generation Layer (see Figure 2).
@@
    • Branding & Asset Panel: Enables users to specify color palettes, typography, and brand values, as well as upload brand assets (logos, images, key visuals).
@@
    • Summary: A high-level overview of the ad concept.
@@ 4.2. Implementation walkthrough
Emily begins by interacting with the Branding Panel, where she inputs key brand attributes that will shape the advertisement. She selects her brand’s colors in the color palette (Figure  3A), specifies Montserrat and Barlow as the preferred typefaces (Figure  3B), and uncertain about a preferred visual aesthetic (e.g., retro, natural, or minimalist), Emily chooses not to define a fixed Ad Style (Figure 3C). As a result, ACAI does not apply any stylistic constraints during initial Ad generation; however, the system remains flexible, allowing Emily to retroactively introduce stylistic preferences at any stage in the co-creation process. This design choice supports customization, even for users who may not have a fully articulated creative vision at the outset. Next, Emily proceeds to the Brand Values (Figure  3D). While she has the option to input a brand description manually, ACAI automatically generates an initial summary after she uploads brand assets (Figure  3E) and selects a target audience segment (Figure  3H). This autogenerated text synthesizes cues from her selected visuals and audience preferences, offering a system-generated interpretation of brand value. The text remains editable both before and after ad generation, enabling iterative refinement. This mechanism supports co-creation by lowering the initial barrier of articulating a brand narrative, particularly for novice users or those with limited experience expressing brand values in textual form. Emily then uploads and selects brand assets (Figure  3E), including Arrivl’s logo and three images from previous advertisements, although other types of visual material may also be used. While ACAI allows users to select between zero and four assets, Emily opts to include all four. These selected assets are treated as priority inputs and prominently featured in the generated ad brief, ensuring that user-provided content substantively informs the system’s output.
With Arrivl’s core branding traits set, Emily moves onto Audience & Goals Panel. Given Arrivl’s focus on travel, Emily enters ‘curious’ and ‘joyful’ as the expected user emotions (Figure 3F). Next, a range of advertisement goals (Figure 3G) are available for selection as well as a text field for a custom goal. Emily selects ‘brand awareness’ as the goal for both Arrivl’s primary and secondary audiences (Figure 3H). To explore potential market segments, Emily engages with the Target Audience (Figure 3H), inputting “Adventurer” and “The Mindful Traveler” as her primary and secondary segments. Based on her earlier inputs, including brand values and selected assets, ACAI generates associated demographic traits, interests, and motivations. This feature supports novice users in articulating audience strategy and allows for further refinement.
@@ 5.1. Reimagining generative workflows
ACAI reimagines how novice users interact with generative AI systems by shifting from conventional text-based prompting to a structured, multimodal interaction paradigm. As a GenAI-powered advertisement creation tool, ACAI is designed to support small business owners in creating brand-aligned advertisements through an interface that reduces the cognitive burden of prompt engineering. By offering a more expressive and structured alternative to conventional text-based input paradigms, ACAI surfaces broader design considerations around ease of prompting and interaction flexibility that are central to evolving AI-augmented UI workflows.
While ACAI does not generate user interfaces per se, it contributes to the workshop’s central theme by demonstrating how AI functionality can inform interface architecture, input design, and creative tooling workflows. The system includes three structured input panels focused on branding, audience and goals, and inspiration, along with multimodal input mechanisms such as textual and image-based style selection. The Inspiration Board enables users to select reference images, which are analyzed by the system to extract design elements such as color, texture, and composition. When users hover over specific elements, the system generates descriptive design text, for example, ”rough textured, sun-drenched rock providing a natural platform and grounding element” and dynamically generates context-specific sliders to adjust stylistic and compositional attributes such as visual emphasis, mood, spatial depth, and aesthetic tone. These inputs guide content generation and demonstrate how UI workflows can be reimagined to better support AI-mediated creation.
@@ 5.2. Toward Inclusive AI Interface Design
In presenting this work, we aim to contribute a perspective on how designing AI-powered interfaces for novice creators can inform the development of more inclusive interfaces. We will highlight how ACAI’s structured and multimodal interface offers alternatives to conventional text-based prompting paradigms, and how such approaches can expand the design space for AI-enabled UI development. In doing so, we invite further reflection on how current generative AI design tooling might evolve beyond expert-oriented workflows toward more scaffolded and equitable interaction paradigms, which we believe are central to the themes this workshop aims to explore.
@@ 6. Conclusion and Future Work
In this paper, we presented ACAI (AI Co-Creation for Advertising and Inspiration), a multimodal generative AI prototype that reimagines the prompt interface for novice designers. Drawing on a formative study with small business owners, we identified key challenges in existing generative AI systems: the cognitive burden of prompt engineering and the generation of generic outputs that fail to align with users’ brand vision. ACAI addresses these challenges through a structured, panel-based interface that supports user input across brand identity, audience targeting, and visual style preferences.
Our approach demonstrates how structured interfaces can foreground user-defined context to improve both alignment and promptability in novice workflows. ACAI contributes to ongoing HCI research on co-creative tooling by offering an accessible, context-aware alternative to conventional prompt-based interaction  paradigms. We refer readers to (Karnatak et al., 2025) for a full discussion of the evaluation, which investigates how ACAI supports small business owners in creating brand-aligned advertisements within situated design workflows. Looking ahead, future research could explore how ACAI’s co-creation model might be adapted to other settings that require contextual grounding but involve users with limited design expertise.

### v2

@@ Abstract.
Text-based prompting remains the predominant interaction paradigm in generative AI, yet it often introduces friction for novice users such as small business owners (SBOs), who struggle to articulate creative goals in domain-specific contexts like advertising. Through a formative study with six SBOs in the United Kingdom, we identify three key challenges: difficulties in expressing brand intuition through prompts, limited opportunities for fine-grained adjustment and refinement during and after content generation, and the frequent production of generic content that lacks brand specificity. In response, we present ACAI (AI Co-Creation for Advertising and Inspiration), a multimodal generative AI tool designed to support novice designers by moving beyond traditional prompt interfaces. ACAI features a structured input system composed of three panels: Branding, Audience and Goals, and the Inspiration Board. These inputs allow users to convey brand-relevant context and visual preferences. This work contributes to HCI research on generative systems by showing how structured interfaces can foreground user-defined context, improve alignment, and enhance co-creative control in novice creative workflows.
@@ 1. Introduction
To investigate how current generative systems support SBO creative workflows, we conducted a formative study with six small business owners (SBOs) in Manchester, United Kingdom. Using semi-structured interviews, we examined participants’ advertising practices, experiences with generative AI, and expectations for AI-assisted design. Our findings surfaced three recurring challenges: (1) the cognitive effort required to translate abstract, domain-specific goals into usable prompts; (2) dissatisfaction with generic outputs that failed to reflect brand tone, values, or audience needs; and (3) concerns about losing creative control due to over-automation, signaling the need for systems that enable co-creation and preserve user agency. These themes align with ongoing HCI work on promptability (Morris, 2024) and the “gulf of envisioning” (Subramonyam et al., 2024), which highlight how non-expert users struggle to effectively guide generative systems.
@@
This work makes three primary contributions to HCI literature:
@@ 2. Related Work
We draw on two strands of prior research that shape our approach. The discussion below reviews foundational work in these areas and situates our contribution as an extension of ongoing efforts to design inclusive, context-aware generative systems.
@@ 2.1. Contextualization in Generative and Interactive Systems
However, current text-to-image generation models primarily rely on text-based prompts and offer limited affordances for encoding structured user intent or domain-specific constraints (Lee et al., 2023). While these systems cproduce compelling outputs, they often fall short in supporting intention-aligned co-creation, particularly for novice users in task-specific domains. Motivated by these limitations, we position ACAI to bridge the gap between generic text-to-image systems and context-sensitive creative workflows by foregrounding explicit, user-defined contextualization.
@@ 3.1. Key Findings from Formative Research
We analyzed the transcripts using reflexive thematic analysis. The following three findings emerged:
@@
    Challenges in Articulating Brand Intuition through Prompts: Participants struggled to articulate their brand vision using text-based prompts. While they possessed a strong intuitive understanding of their brand identity, they found it difficult to externalize this understanding. One participant shared, “I instructed it to be funny, but that’s quite subjective, so I had to edit it quite heavily.” –P2. This finding underscores the tacit and iterative nature of branding, highlighting how limitations in promptability introduce friction into the creative workflows of novice users.
@@
    Generic and Misaligned AI Outputs: Participants frequently described AI-generated advertisements as generic, repetitive, or misaligned with their brand’s tone and aesthetic. They noted that outputs often lacked the nuance necessary to reflect distinctive brand identities. For example, one participant reflected: “I was looking at some of these AI branding things, but nothing came of it, what I was offered… it was quite childish, quite gimmicky.” – P1. This disconnect reinforces the need for generative systems that incorporate user-defined brand elements and contextual cues to produce brand-consistent outputs.
@@
    Tensions Between AI-Generated Content and User Control: While participants appreciated the efficiency enabled by generative AI, many voiced concerns about single-prompt execution, where a text prompt yields a output that cannot be directly edited. Instead, users were required to craft new prompts to regenerate what they wanted, often through trial and error. Participants were not opposed to AI-generated content itself; rather, their frustration stemmed from outputs that misaligned with their intent and lacked accessible pathways for refinement. As one participant explained, “You couldn’t really change it. When AI produced that for you, there weren’t tweaks you could make to it… so that’s the limitation of it” –P3. These reflections underscore the importance of designing interaction models that support direct editability, iterative control, and co-creative engagement.
@@ 4. ACAI: AI Co-Creation for Advertising and Inspiration
In response to these requirements, we developed ACAI (AI Co-Creation for Advertising and Inspiration), a generative AI-powered prototype designed to support the creation of brand-aligned advertisements. ACAI is designed for small business owners (SBOs) who possess deep contextual knowledge of their brand but lack formal design expertise. By aligning generative AI capabilities with users’ domain-specific insights, ACAI facilitates a co-creative design process.
@@ 4.1. System Overview
ACAI leverages Gemini 1.5 Pro, a multimodal large language model (MLLM), to support its generative capabilities. ACAI’s architecture is organized into three core components: the User Input Layer, the Prompt Processing Layer, and the Output Generation Layer (see Figure 2).
@@
    • Branding Panel: Enables users to specify color palettes, typography, and brand values, as well as upload brand assets (logos, images, key visuals).
@@
    • Summary: A high-level overview of the advertisement concept.
@@ 4.2. Implementation walkthrough
Emily begins by interacting with the Branding Panel, where she inputs key brand attributes that will shape the advertisement. She selects her brand’s colors in the color palette (Figure  3A), specifies Montserrat and Barlow as the preferred typefaces (Figure  3B), and uncertain about a preferred visual aesthetic (e.g., retro, natural, or minimalist), Emily chooses not to define a fixed Ad Style (Figure 3C). As a result, ACAI does not apply any stylistic constraints during initial advertisement generation; however, the system remains flexible, allowing Emily to retroactively introduce stylistic preferences at any stage in the co-creation process. This design choice supports customization, even for users who may not have a fully articulated creative vision at the outset. Next, Emily proceeds to the Brand Values (Figure  3D). While she has the option to input a brand description manually, ACAI automatically generates an initial summary after she uploads brand assets (Figure  3E) and selects a target audience segment (Figure  3H). This autogenerated text synthesizes cues from her selected visuals and audience preferences, offering a system-generated interpretation of brand value. The text remains editable both before and after advertisement generation, enabling iterative refinement. This mechanism supports co-creation by lowering the initial barrier of articulating a brand narrative, particularly for novice users or those with limited experience expressing brand values in textual form. Emily then uploads and selects brand assets (Figure  3E), including Arrivl’s logo and three images from previous advertisements, although other types of visual material may also be used. While ACAI allows users to select between zero and four assets, Emily opts to include all four. These selected assets are treated as priority inputs and prominently featured in the generated Ad brief, ensuring that user-provided content substantively informs the system’s output.
With Arrivl’s core branding traits set, Emily moves onto Audience & Goals Panel. Given Arrivl’s focus on travel, Emily enters ‘curious’ and ‘joyful’ as the expected user emotions (Figure 3F). Next, a range of advertisement goals (Figure 3G) are available for selection as well as a text field for a custom goal. Emily selects ‘brand awareness’ as the goal for both Arrivl’s primary and secondary audiences (Figure 3H). To explore potential market segments, Emily interacts with the Target Audience panel (Figure 3H), entering “Adventurer” and “The Mindful Traveler” as her primary and secondary segments. Drawing on these entries, along with information from other ACAI panels that is either manually input by Emily or automatically generated, ACAI produces corresponding demographic traits, interests, and motivations. This feature is designed to help novice users articulate their audience strategy and offers opportunities for further refinement.
@@ 5.1. Reimagining generative workflows
ACAI reimagines how novice users engage with generative AI by moving beyond conventional free-form text prompting toward a structured, multimodal interaction paradigm. Unlike systems that rely primarily on open-ended textual input, ACAI enables users to steer content generation through clearly defined input panels and diverse input modalities. It is designed to help small business owners create brand-aligned content through an interface that reduces the cognitive effort required for prompt formulation.
By offering a more expressive and guided alternative to traditional prompting, ACAI highlights broader design implications related to interaction flexibility, editability, and promptability. These are critical considerations in the development of evolving AI-augmented UI workflows. While ACAI does not generate UI layouts directly, it contributes to the workshop’s central theme by demonstrating how AI functionality can inform interface architecture, input design, and creative tooling practices.
ACAI’s interface, described in detail in Section 4.2, illustrates how structured input panels, multimodal selection mechanisms, visual style extraction, and dynamically generated sliders enable users to adjust visual attributes such as mood, spatial depth, and composition. These affordances reflect a broader reimagining of how generative systems can support control, co-creative flexibility, and intent alignment within creative workflows.
@@ 5.2. Toward Inclusive AI Interface Design
Through this work, we offer a perspective on how designing AI-powered interfaces for novice creators can inform the development of more inclusive interaction paradigms. We highlight how ACAI’s structured and multimodal interface provides alternatives to dominant text-based prompting approaches, thereby expanding the design space for generative AI user interface development. In articulating these possibilities, we aim to provoke broader reflection on how generative AI tooling might move beyond expert-centric workflows toward more scaffolded and equitable interaction paradigms.
@@ 6. Conclusion and Future Work
In this paper, we introduced ACAI (AI Co-Creation for Advertising and Inspiration), a multimodal generative AI prototype designed to support novice designers through a structured alternative to traditional prompt-based interfaces. Our formative study with small business owners revealed three persistent limitations in current generative workflows: difficulty in externalizing brand intuition through text, limited opportunities for fine-grained control during and after content generation, and the frequent production of generic outputs misaligned with brand identity. ACAI addresses these challenges through a panel-based interface that enables users to specify branding elements, define audience goals, and select visual inspirations, facilitating more directed and brand-consistent co-creation. Our approach demonstrates how structured interfaces can foreground user-defined context to improve both alignment and promptability in novice workflows. ACAI contributes to ongoing HCI research on co-creative tooling by offering an accessible, context-aware alternative to conventional prompt-based interaction  paradigms. We refer readers to (Karnatak et al., 2025) for a full discussion of the evaluation, which investigates how ACAI supports small business owners in creating brand-aligned advertisements within situated design workflows. Looking ahead, future research could explore how ACAI’s co-creation model might be adapted to other settings that require contextual grounding but involve users with limited design expertise.